OpenStudy (anonymous):

What is the value of the variable that makes the statement true? b + (–2) + 8 = 0 A. –10 B. –6 C. 6 D. 10

OpenStudy (igreen):

It wants you to solve for b. b + (–2) + 8 = 0 b - 2 + 8 = 0 Add: b + 6 = 0 Subtract 6 to both sides: b = -6

OpenStudy (igreen):

So your answer is B..

OpenStudy (ahsome):

^^ Yes ^^